(Q)Consider a one-dimensional box extends from x = x0 to x = x0 + L, where x0 6= 0.




Would the expression for the wave function of a particle in this box be any different




from those in the box that extends from x = 0 to x = L? Would the energy levels be




different?





1
Expert's answer
2022-02-06T14:34:07-0500

The expression for the wavefunction for the box that  extends from x = 0 to x = L is given as follows (see https://chem.libretexts.org/Bookshelves/Physical_and_Theoretical_Chemistry_Textbook_Maps/Supplemental_Modules_(Physical_and_Theoretical_Chemistry)/Quantum_Mechanics/05.5%3A_Particle_in_Boxes/Particle_in_a_1-Dimensional_box):


"\psi=\sqrt{\dfrac{2}{L}}\sin\dfrac{n\pi}{L}x"

Since this expression depends on "x", then the coordinate shift "x\to x-x_0" wich correspons to the  one-dimensional box extends from x = x0 to x = x0 + L will lead to the following expression:


"\psi=\sqrt{\dfrac{2}{L}}\sin\dfrac{n\pi}{L}(x-x_0)"

Thus, the wavefunction has changed.


On the other hand, the expression for energy levels depends only on "L" (the width of the box):


"E_n = \dfrac{n^2h^2}{8mL^2}"

thus, remains unchanged under the box shift.